OpenStudy (anonymous):

Simplify 12A + 2 + A - 1.

pooja195 (pooja195):

Combine like terms

OpenStudy (anonymous):

16x?

OpenStudy (haseeb96):

just add those same terms like 12A +A +2 -1 13A +1 is your answer

pooja195 (pooja195):

Dogs+Dogs= ? Dogs = 12A+A Cat+cats=? Cats=2-1

OpenStudy (anonymous):

?

pooja195 (pooja195):

12+1=?

OpenStudy (anonymous):

13

pooja195 (pooja195):

So we have 13A 2-1=?

OpenStudy (anonymous):

1

pooja195 (pooja195):

13A+1 would be what you are left with : )

pooja195 (pooja195):

And thats our final answer

OpenStudy (anonymous):

oh okay thank you

pooja195 (pooja195):

:)